The probability that it will rain on any given day is x. If it is raining, the probability that Aran wears a hat is 0.8 and if it is not raining, the probability that he wears a hat is 0.3. Whether it is raining or not, if Aran wears a hat, the probability that he wears a scarf is 0.4. If he does not wear a hat, the probability that he wears a scarf is 0.1. The probability that on a randomly chosen day it is not raining and Aran is not wearing a hat or a scarf is 0.36. Find the value of x.
Mark scheme
Marking points from the official mark scheme (Feb/Mar 2023). The mark or mark type is shown in brackets.
- [M1] (1-x)×a×b=0.36, a=0.7 or 0.3, b=0.9or0.1
- [B1] (1-x)×0.7×0.9=0.36, (1-x)×0.63=0.36, 0.63 - 0.63x = 0.36 or 1-x=0.36/0.63 seen. Condone recovery from omission of brackets.
- [A1] x=3/7. Accept 0.428571 to at least 3 sf. Condone 0.4285 rounding to 0.429. If M0 awarded, SC B1 for x = 3/7 or 0.428571 to at least 3 sf.
Write your answer and get it marked
On Oracle Prep you can write or photograph your own answer and have it marked by AI against this mark scheme, with feedback on what you missed.
Sign up free →